### v2

@@ Abstract
Large Language Models (LLMs) are revolutionizing the development of AI assistants capable of performing diverse tasks across domains. However, current state-of-the-art LLM-driven agents face significant challenges, including high operational costs and limited success rates on complex benchmarks like GAIA. To address these issues, we propose the Knowledge Graph of Thoughts (KGoT), an innovative AI assistant architecture that integrates LLM reasoning with dynamically constructed knowledge graphs (KGs). KGoT extracts and structures task-relevant knowledge into a dynamic KG representation, iteratively enhanced through external tools such as math solvers, web crawlers, and Python scripts. Such structured representation of task-relevant knowledge enables low-cost models to solve complex tasks effectively. For example, KGoT achieves a 29% improvement in task success rates on the GAIA benchmark compared to Hugging Face Agents with GPT-4o mini, while reducing costs by over 36$\times$ compared to GPT-4o. Improvements for recent reasoning models are similar, e.g., 36% and 37.5% for Qwen2.5-32B and Deepseek-R1-70B, respectively. KGoT offers a scalable, affordable, and high-performing solution for AI assistants.
@@ 1 Introduction
Large Language Models (LLMs) are transforming the world. However, training LLMs is expensive, time-consuming, and resource-intensive. In order to democratize the access to generative AI, the landscape of agent systems has massively evolved during the last two years LangChain Inc. (2024); Rush (2023); Kim et al. (2024); Sumers et al. (2024); Hong et al. (2024); Guo et al. (2024); Edge et al. (2024); Besta et al. (2025b); Zhuge et al. (2024); Beurer-Kellner et al. (2024); Shinn et al. (2023); Kagaya et al. (2024); Zhao et al. (2024a); Stengel-Eskin et al. (2024); Wu et al. (2024). These schemes have been applied to numerous tasks in reasoning Creswell et al. (2023); Bhattacharjya et al. (2024); Besta et al. (2025b), planning Wang et al. (2023c); Prasad et al. (2024); Shen et al. (2023); Huang et al. (2023), software development Tang et al. (2024), and many others Xie et al. (2024); Schick et al. (2023); Beurer-Kellner et al. (2023).
@@
The GAIA benchmark Mialon et al. (2024) has emerged as a valuable standard for evaluating LLM-driven agent architectures in their capacity to function as general-purpose AI assistants. This benchmark rigorously tests these systems across diverse tasks (involving web navigation, code execution, image reasoning, scientific QA, and multimodal tasks), providing a clear measure of their competence. However, despite more than a year since its introduction, the top-performing solutions on GAIA still fail at many tasks. Furthermore, the cost of operating these systems is prohibitively high. For instance, executing all tasks from the validation set using Hugging Face Agents Roucher & Petrov (2024) with GPT-4o incurs costs of roughly \$200, illustrating the need for more cost-efficient alternatives. While deploying smaller models like GPT-4o mini offers significant cost reductions, it results in a substantial decline in task success rates, rendering it an insufficient solution. Alternatively, when using open models, maintaining the infrastructure for large models is costly and often prohibitive for an average user; small open models require inexpensive hardware but are less capable.
Figure 1: Illustration of the key idea behind Knowledge Graph of Thoughts (KGoT): transforming the representation of a task for an AI assistant from a textual form into a knowledge graph (KG). As an example, we use a Level-3 (i.e., highest difficulty) task from the GAIA benchmark. In order to solve the task, KGoT evolves this KG by adding relevant information that brings the task closer to completion. This is achieved by iteratively running various tools. Finally, the task is solved by extracting the relevant information from the KG, using – for example – a graph query, or an LLM’s inference process with the KG provided as a part of the input prompt.
To address these challenges, we propose the Knowledge Graph of Thoughts (KGoT), a novel AI assistant architecture designed to significantly reduce task execution costs while maintaining a high success rate (contribution #1). The central innovation of KGoT lies in its use of a knowledge graph (KG) Singhal (2012); Besta et al. (2024b) to extract and structure knowledge relevant to a given task. A KG organizes information into triples, providing a structured representation of knowledge that small, cost-effective models can efficiently process. Hence, KGoT “turns the unstructured into the structured”, i.e., KGoT turns the often unstructured data such as website contents or PDF files into structured KG triples. This approach enhances the comprehension of task requirements, enabling even smaller models to achieve performance levels comparable to much larger counterparts, but at a fraction of the cost.
@@
Our evaluation against top GAIA leaderboard baselines demonstrates its effectiveness and efficiency (contribution #3). KGoT solves $>$2$\times$ more tasks from the validation set than Hugging Face Agents with GPT-4o mini. Moreover, harnessing a smaller model dramatically reduces operational costs. Specifically, with GPT-4o mini instead of GPT-4o, KGoT lowers task execution costs from \$187 to roughly \$5. On top of that, KGoT reduces bias and improves fairness by externalizing reasoning into an explicit knowledge graph rather than relying solely on the LLM’s internal generation. This ensures that key steps when resolving tasks are grounded in transparent and auditable information. This highlight the potential of KGoT in developing affordable AI assistants capable of high performance across a diverse range of tasks.
@@ 2.1 What is a Knowledge Graph?
Mathematically, a knowledge graph can be defined as a directed labeled graph $G=(V,E,L)$, where $V$ is the set of vertices (entities), $E\subseteq V\times V$ is the set of edges (relationships), and $L$ is the set of labels (predicates) assigned to the edges. Each entity or predicate may further include properties or attributes, enabling richer representation. Knowledge graphs are widely used in various domains, including search engines, recommendation systems, and AI reasoning, as they facilitate both efficient storage and complex queries.
2.2 Harnessing KGs for Effective Task Resolution
@@
At the heart of KGoT is the process of transforming a task solution state into an evolving KG. The KG representation of the task is built from “thoughts” generated by the LLM in the iterative process of enhancing the KG. These “thoughts” are intermediate insights identified by the LLM as it works through the problem. Each thought contributes to refining or expanding the KG by adding new vertices, edges, or attributes that represent new information.
@@
For example, consider the following Level 3 (i.e., highest difficulty) task from the GAIA benchmark: “In the YouTube 360 VR video from March 2018 narrated by the voice actor of Lord of the Rings’ Gollum, what number was mentioned by the narrator directly after dinosaurs were first shown in the video?” (see Figure 1). Here, the KG representation of the task solution state has a vertex “Gollum (LotR)”. Then, the thought “Gollum from Lord of the Rings is interpreted by Andy Serkis” results in adding a vertex for “Andy Serkis”, and linking “Gollum (LotR)” to “Andy Serkis” with the predicate “interpreted by”. Such integration of thought generation and KG construction creates a feedback loop where the KG continuously evolves as the task progresses, aligning the representation with problem requirements.
@@
In order to evolve the KG task representation, KGoT interacts with tools and retrieves more information. For instance, the system might query the internet to identify videos narrated by Andy Serkis (e.g., “The Silmarillion“ and “We Are Stars”). It can also use a YouTube transcriber tool to find their publication date, adding new vertices and edges to the graph. Intermediate results, such as the video type or comparisons, are incorporated back into the graph, creating a more complete and structured representation of the task.
This iterative refinement allows the KG to model the current “state” of the task at each step, bringing it closer to completion. The system’s dynamic nature enables it to address a wide range of tasks by adapting the graph’s structure and content in response to real-time interactions. For example, multi-step reasoning tasks, such as synthesizing data from different sources or performing calculations, are handled by adding relevant subgraphs or updating existing vertices based on the latest retrieved information. Once the KG has been sufficiently populated with task-specific knowledge, it serves as a robust resource for solving the problem.
2.3 Extracting Information from KG
To accommodate varying performance requirements and tasks, KGoT supports different ways to extract the information from the KG when solving a task. Currently, we offer graph query languages or general-purpose languages; each of them can be combined with the so-called Direct Retrieval.
Graph Query Languages First, to solve the task, one can use a graph query, prepared by the LLM in a language such as Cypher Francis et al. (2018) or SPARQL Pérez et al. (2009), to extract the answer to the task from the graph. This capability is particularly advantageous for tasks that require retrieving specific subgraphs, relationships, or patterns within the KG.
@@ 2.3 Extracting Information from KG
General-Purpose Languages Another way to extract the information from the KG that is needed to solve the task is to use a script prepared by the LLM in a general-purpose programming language such as Python. This approach, while not as effective as query languages for workloads such as pattern matching, offers greater flexibility and may outperform the latter when a task requires, for example, traversing a long path in the graph.
Direct Retrieval In certain cases, once enough information is gathered into the KG, it may be more effective to directly paste the KG into the LLM context and ask the LLM to solve the task, instead of preparing a dedicated query or script. We refer to this approach as Direct Retrieval.
@@
Accuracy-Cost-Runtime Tradeoff The three above schemes offer a tradeoff between accuracy, cost, and runtime. For example, when low latency is of top priority, general-purpose languages and the corresponding frameworks such as NetworkX should be used, as they provide an efficient lightweight representation of the KG and offer rapid access and modification of graph data. When token cost is most important, one should avoid Direct Retrieval (which consumes many tokens as it directly embeds the KG into the LLM context) and focus on either query or general-purpose languages, with a certain preference for the former, because – based on our experience – generated queries tend to be shorter than scripts. Finally, when aiming for solving as many tasks as possible, one should experiment with all three schemes –As shown in the Evaluation section, these methods have complementary strengths—Direct Retrieval is effective for broad contextual understanding, while graph queries and scripts are better suited for structured reasoning.
@@ 3 System Architecture
The KGoT system, pictured in Figure 2, is designed as a modular and flexible framework that comprises three main components: the Graph Store Module, the Controller, and the Integrated Tools, each playing a critical role in the task-solving process. Below, we provide a detailed description of each component and its role in the system.
3.1 Graph Store Module for Maintaining the KG
A key component of the KGoT system is the Graph Store Module, which manages the storage and retrieval of the dynamically evolving knowledge graph which represents the task state. In order to harness graph queries, we use a graph database backend; in the current KGoT implementation, it is Neo4j Robinson et al. (2015), an established graph database Besta et al. (2023c; b; d) (we selected Cypher and Neo4j after an analysis of the literature that indicated limitations for the LLM-based query generation of SPARQL queries Emonet et al. (2025); Mecharnia & d’Aquin (2025)). Then, in order to support graph accesses using a general-purpose language, KGoT harnesses the NetworkX library NetworkX Developers (2024) and Python. Note that the extensible design of KGoT enables seamless integration of any other backends and languages.
3.2 Controller for Workflow Management
The Controller is the central orchestrator of the KGoT system, responsible for managing the interaction between the knowledge graph and the integrated tools. When a user submits a query, the Controller initiates the reasoning process by interpreting the task and coordinating the steps required for resolution. It dynamically determines which tools to invoke based on the current state of the KG and the specific requirements of the task. As tools produce results, the Controller integrates these outputs back into the KG, updating its structure to reflect the new knowledge.
The KGoT Controller employs a dual-LLM architecture with a clear separation of roles between constructing the KG (managed by the LLM Graph Executor) and interacting with tools (managed by the LLM Tool Executor).
The LLM Graph Executor determines the next steps after each iteration that constructs and evolves the KG. It identifies any missing information necessary to solve the task, formulates appropriate queries for interacting with the graph store backend (retrieve/insert operations), and parses intermediate or final results for integration into the KG. It also prepares the final response to the user by synthesizing outputs from all previous steps and from the KG.
@@ 3.2 Controller for Workflow Management
3.3 Integrated Tools for Evolving KG
The Integrated Tools Module in the KGoT system provides a hierarchical and diverse suite of specialized tools, each tailored to address specific task requirements. At the foundation is the Python Code Tool, enabling the generation and execution of dynamic scripts for complex computations and algorithmic tasks. The code tool is also used when solving math steps. Supplementing the controller’s reasoning, the LLM Tool integrates an additional language model to provide extended knowledge beyond the constrained capabilities of the controller’s LLM, ensuring robust reasoning with minimal risk of hallucinations. For multimodal tasks, the Image Tool facilitates image processing and information extraction. Web-based operations are handled by the Surfer Agent, based on the design by Hugging Face Agents Roucher & Petrov (2024), which interacts with the web through tools like the Wikipedia Tool and granular navigation tools (e.g., PageUp, PageDown, Find) while leveraging SerpApi SerpApi LLM (2025) for effective searches. Additional capabilities include the ExtractZip Tool, designed for processing compressed archives, and the Text Inspector Tool, which extracts and transforms text from diverse sources such as MP3 files, YouTube transcripts, and various formats into Markdown. This modular hierarchy ensures flexibility, extensibility, and adaptability in solving a wide range of complex tasks.
3.4 High-Performance & Scalability
@@
KGoT uses various optimizations to enhance scalability and performance. They include (1) asynchronous execution using asyncio Python Software Foundation (2025b) to parallelize LLM tool invocations, mitigating I/O bottlenecks and reducing idle time, (2) graph operation parallelism by reformulating LLM-generated Cypher queries to enable concurrent execution of independent operations in a graph database, and (3) MPI-based distributed processing, which decomposes workloads into atomic tasks distributed across ranks using a work-stealing algorithm to ensure balanced computational load and scalability.
3.5 System Robustness with Majority Voting

### v6

@@ Abstract
[…]
@@ 1 Introduction
Large Language Models (LLMs) are transforming the world. However, training LLMs is expensive, time-consuming, and resource-intensive. In order to democratize the access to generative AI, the landscape of agent systems has massively evolved during the last two years (LangChain Inc., 2025a; Rush, 2023; Kim et al., 2024; Sumers et al., 2024; Hong et al., 2024; Guo et al., 2024; Edge et al., 2025; Besta et al., 2025c; Zhuge et al., 2024; Beurer-Kellner et al., 2024; Shinn et al., 2023; Kagaya et al., 2024; Zhao et al., 2024a; Stengel-Eskin et al., 2024; Wu et al., 2024). These schemes have been applied to numerous tasks in reasoning (Creswell et al., 2023; Bhattacharjya et al., 2024; Besta et al., 2025c), planning (Wang et al., 2023c; Prasad et al., 2024; Shen et al., 2023; Huang et al., 2023), software development (Tang et al., 2024), and many others (Xie et al., 2024; Li & Vasarhelyi, 2024; Schick et al., 2023; Beurer-Kellner et al., 2023).
@@
The GAIA benchmark (Mialon et al., 2024) has become a key standard for evaluating LLM-based agent systems across diverse tasks, including web navigation, code execution, image reasoning, scientific QA, and multimodal challenges. Despite its introduction nearly two years ago, top-performing solutions still struggle with many tasks. Moreover, operational costs remain high: running all validation tasks with Hugging Face Agents (Roucher & Petrov, 2025) and GPT-4o costs $\approx$\$200, underscoring the need for more affordable alternatives . Smaller models like GPT-4o mini significantly reduce expenses but suffer from steep drops in task success, making them insufficient. Open large models also pose challenges due to demanding infrastructure needs, while smaller open models, though cheaper to run, lack sufficient capabilities.
To address these challenges, we propose Knowledge Graph of Thoughts (KGoT), a novel AI assistant architecture that significantly reduces task execution costs while maintaining a high success rate (contribution #1). The central innovation of KGoT lies in its use of a knowledge graph (KG) (Singhal, 2012; Besta et al., 2024b) to represent knowledge relevant to a given task. A KG organizes information into triples, providing a structured representation of knowledge that small, cost-effective models can efficiently process. Hence, KGoT “turns the unstructured into the structured”, i.e., KGoT turns the often unstructured data such as website contents or PDF files into structured KG triples. This approach enhances the comprehension of task requirements, enabling even smaller models to achieve performance levels comparable to much larger counterparts, but at a fraction of the cost.
@@
[…]
On top of that, KGoT reduces noise and simultaneously minimizes bias and improves fairness by externalizing reasoning into an explicit knowledge graph rather than relying solely on the LLM’s internal generation (contribution #4). This ensures that key steps when resolving tasks are grounded in transparent, explainable, and auditable information.
@@ 2.1 What is a Knowledge Graph?
A knowledge graph (KG) is a structured representation of information that organizes knowledge into a graph-based format, allowing for efficient querying, reasoning, and retrieval. Formally, a KG consists of a set of triples, where each triple $(s,p,o)$ represents a relationship between two entities $s$ (subject) and $o$ (object) through a predicate $p$. For example, the triple $(\text{``Earth''},\text{``orbits''},\text{``Sun''})$ captures the fact that Earth orbits the Sun. Mathematically, a knowledge graph can be defined as a directed labeled graph $G=(V,E,L)$, where $V$ is the set of vertices (entities), $E\subseteq V\times V$ is the set of edges (relationships), and $L$ is the set of labels (predicates) assigned to the edges. Each entity or predicate may further include properties or attributes, enabling richer representation. Knowledge graphs are widely used in various domains, including search engines, recommendation systems, and AI reasoning, as they facilitate both efficient storage and complex queries.
Figure 1: The key idea behind Knowledge Graph of Thoughts (KGoT): transforming the representation of a task for an AI assistant from a textual form into a knowledge graph (KG). As an example, we use a Level-3 (i.e., highest difficulty) task from the GAIA benchmark. In order to solve the task, KGoT evolves this KG by adding relevant information that brings the task closer to completion. This is achieved by iteratively running various tools. Finally, the task is solved by extracting the relevant information from the KG, using – for example – a graph query, or an LLM’s inference process with the KG provided as a part of the input prompt. More examples of KGs are in Appendix A.
2.2 Harnessing Knowledge Graphs for Effective AI Assistant Task Resolution
@@
At the heart of KGoT is the process of transforming a task solution state into an evolving KG. The KG representation of the task is built from “thoughts” generated by the LLM. These “thoughts” are intermediate insights identified by the LLM as it works through the problem. Each thought contributes to expanding or refining the KG by adding vertices or edges that represent new information.
@@
For example, consider the following Level 3 (i.e., highest difficulty) task from the GAIA benchmark: “In the YouTube 360 VR video from March 2018 narrated by the voice actor of Lord of the Rings’ Gollum, what number was mentioned by the narrator directly after dinosaurs were first shown in the video?” (see Figure 1 for an overview; more examples of constructed KGs are in Appendix A). Here, the KG representation of the task solution state has a vertex “Gollum (LotR)”. Then, the thought “Gollum from Lord of the Rings is interpreted by Andy Serkis” results in adding a vertex for “Andy Serkis”, and linking “Gollum (LotR)” to “Andy Serkis” with the predicate “interpreted by”. Such integration of thought generation and KG construction creates a feedback loop where the KG continuously evolves as the task progresses, aligning the representation with problem requirements.
@@
In order to evolve the KG task representation, KGoT iteratively interacts with tools and retrieves more information. For instance, the system might query the internet to identify videos narrated by Andy Serkis (e.g., “The Silmarillion“ and “We Are Stars”). It can also use a YouTube transcriber tool to find their publication date. This iterative refinement allows the KG to model the current “state” of a task at each step, creating a more complete and structured representation of this task and bringing it closer to completion. Once the KG has been sufficiently populated with task-specific knowledge, it serves as a robust resource for solving the problem.
In addition to adding new graph elements, KGoT also supports other graph operations. This includes removing nodes and edges, used as a part of noise elimination strategies.
2.3 Extracting Information from the KG
@@
To accommodate different tasks, KGoT supports different ways to extract the information from the KG. Currently, we offer graph query languages or general-purpose languages; each of them can be combined with the so-called Direct Retrieval. First, one can use a graph query, prepared by the LLM in a language such as Cypher (Francis et al., 2018) or SPARQL (Pérez et al., 2009), to extract the answer to the task from the graph. This works particularly well for tasks that require retrieving specific patterns within the KG. Second, we also support general scripts prepared by the LLM in a general-purpose programming language such as Python. This approach, while not as effective as query languages for pattern matching, offers greater flexibility and may outperform the latter when a task requires, for example, traversing a long path in the graph. Third, in certain cases, once enough information is gathered into the KG, it may be more effective to directly paste the KG into the LLM context and ask the LLM to solve the task, instead of preparing a dedicated query or script. We refer to this approach as Direct Retrieval.
@@
The above schemes offer a tradeoff between accuracy, cost, and runtime. For example, when low latency is priority, general-purpose languages should be used, as they provide an efficient lightweight representation of the KG and offer rapid access and modification of graph data. When token cost is most important, one should avoid Direct Retrieval (which consumes many tokens as it directly embeds the KG into the LLM context) and focus on either query or general-purpose languages, with a certain preference for the former, because its generated queries tend to be shorter than scripts. Finally, when aiming for solving as many tasks as possible, one should experiment with all three schemes. As shown in the Evaluation section, these methods have complementary strengths: Direct Retrieval is effective for broad contextual understanding, while graph queries and scripts are better suited for structured reasoning.
2.4 Representing the KG
KGoT can construct three interoperable KG representations: Property graphs (used with graph query languages such as Cypher and systems such as Neo4j (Robinson et al., 2015)), RDF graphs (used with graph query languages such as SPARQL and systems such as RDF4J (Ben Mahria et al., 2021)), and the adjacency list graphs (Besta et al., 2018) (used with general-purpose languages such as Python and systems such as NetworkX (NetworkX Developers, 2025)).
Each representation supports a different class of analysis. The Property graph view facilitates analytics such as pattern matching, filtering, of motif queries directly on the evolving task-state graph. The RDF graph view facilitates reasoning over ontology constraints, schema validation, and SPARQL-based inference for missing links. The adjacency list representation with NetworkX facilitates Python-based graph analytics, for example centrality measures, connected components, clustering coefficients, etc., all on the same KG snapshots.
Appendix A contains examples of task-specific KGs, illustrating how their topology varies with the task domain (e.g., tree-like procedural chains vs. dense relational subgraphs in multi-entity reasoning).
2.5 Bias, Fairness, and Noise Mitigation through KG-Based Representation
KGoT externalizes and structures the reasoning process, which reduces noise, mitigates model bias, and improves fairness, because in each iteration both the outputs from tools and LLM thoughts are converted into triples and stored explicitly. Unlike opaque monolithic LLM generations, this fosters transparency and facilitates identifying biased inference steps. It also facilitates noise mitigation: new triples can be explicitly checked for the quality of their information content before being integrated into the KG, and existing triples can also be removed if they are deemed redundant (examples of such triples that have been found and removed are in Appendix B.6).
@@ 3 System Architecture
The KGoT modular and flexible architecture, pictured in Figure 2, consists of three main components: the Graph Store Module, the Controller, and the Integrated Tools, each playing a critical role in the task-solving process. Below, we provide a detailed description of each component and its role in the system. Additional details are in Appendix B (architecture) and in Appendix C (prompts).
3.1 Maintaining the Knowledge Graph with the Graph Store Module
A key component of the KGoT system is the Graph Store Module, which manages the storage and retrieval of the dynamically evolving knowledge graph which represents the task state. In order to harness graph queries, we use a graph database backend; in the current KGoT implementation, we test Cypher together with Neo4j (Robinson et al., 2015), an established graph database (Besta et al., 2023b; c), as well as SPARQL together with the RDF4J backend (Ben Mahria et al., 2021). Then, in order to support graph accesses using a general-purpose language, KGoT harnesses the NetworkX library (NetworkX Developers, 2025) and Python. Note that the extensible design of KGoT enables seamless integration of any other backends and languages.
3.2 Managing the Workflow with the Controller Module
The Controller orchestrates the interactions between the KG and the tools. Upon receiving a user query, it iteratively interprets the task, determines the appropriate tools to invoke based on the KG state and task needs, and integrates tool outputs back into the KG. The Controller uses a dual-LLM architecture with a clear separation of roles: the LLM Graph Executor constructs and evolves the KG, while the LLM Tool Executor manages tool selection and execution.
The LLM Graph Executor determines the next steps after each iteration that constructs and evolves the KG. It identifies any missing information necessary to solve the task, formulates appropriate queries for the graph store interaction (retrieve/insert operations), and parses intermediate or final results for integration into the KG. It also prepares the final response to the user based on the KG.
@@ 3.2 Managing the Workflow with the Controller Module
3.3 Ensuring Versatile and Extensible Set of Integrated Tools
KGoT offers a hierarchical suite of tools tailored to diverse task needs. The Python Code Tool enables dynamic script generation and execution for complex computations. The LLM Tool supplements the controller’s reasoning by integrating an auxiliary language model, enhancing knowledge access while minimizing hallucination risk. For multimodal inputs, the Image Tool supports image processing and extraction. Web-based tasks are handled by the Surfer Agent (based on the design by Hugging Face Agents (Roucher & Petrov, 2025)), which leverages tools like the Wikipedia Tool, granular navigation tools (PageUp, PageDown, Find), and SerpApi (SerpApi LLM, 2025) for search. Additional tools include the ExtractZip Tool for compressed files and the Text Inspector Tool for converting content from sources like MP3s and YouTube transcripts into Markdown. Finally, the user can seamlessly add a new tool by initializing the tool, passing in the logger object for tool use statistics, and appending the tool to the tool list via a Tool Manager object. We require all tools implemented to adhere to the LangChain’s BaseTool interface class. This way, the list of tools managed by the Tool Manager can be directly bound to the LLM Tool Executor via LangChain bind_tools, further facilitating new tools.
3.4 Ensuring High-Performance & Scalability
@@
The used scalability optimizations include (1) asynchronous execution using asyncio (Python Software Foundation, 2025b) to parallelize LLM tool invocations, mitigating I/O bottlenecks and reducing idle time, (2) graph operation parallelism by reformulating LLM-generated Cypher queries to enable concurrent execution of independent operations in a graph database, and (3) MPI-based distributed processing, which decomposes workloads into atomic tasks distributed across ranks using a work-stealing algorithm to ensure balanced computational load and scalability.
3.5 Ensuring System Robustness